# Learning from Emergence: A Study on Proactively Inhibiting the Monosemantic Neurons of Artificial Neural Networks

### v2

@@ E.1. Configuration of Base Models
For the language task, we insert 4 MEemL layers after the 4 middle Attention layers of BERT-base (12 in total). Since MEemL does not introduce any additional parameters, we can directly use their checkpoint after pretraining to finetune. The finetuning setting is the same as BERT for fairness, where each task is trained with a batch size of 32 and 3 epochs over the data for all GLUE tasks. For each task, we selected the best fine-tuning learning rate ($l\in\{5e-5,4e-5,3e-5,2e-5\}$) on the validation dataset (DEV). Baseline results are from Table.1 in BERT (Devlin et al., 2019).
For the image classification task, we also insert 4 MEemL layers after the 4 middle Attention layers for each model. The training is conducted on ImageNet-1k for 30 epochs using pretrained checkpoints on ImageNet-22k. The settings are directly adopted from their scripts, except for the number of GPU cards. Baseline results are collected by running checkpoints from the github of Swin-Transformer with some updates.
For the precipitation task, we insert 6 MEemL lay- ers after the hidden states of each ConvGRU module. The training is conducted on the HKO-7 dataset (Shi et al., 2017) of radar images with a granularity of $120\times 120$, which are centered in Hong Kong. These radar images cover a timespan from 2009 to 2015 and are captured every 6 minutes. The original rainfall intensity is mapped to a scale of [0, 255], forming a regression task. The pretraining settings are the same with theirs, optimizing for 20k steps. Based on that, we conduct 2k more steps funetuning to obtain the final output.
@@ E.3. Parameter Searching over Inhibition Level of Monosemanticity
Table 5. Validation results on GLUE Test datasets. The settings are the same with Table. 1. All metrics are the larger the better with best results in bold font.
Model MNLI-(M/MM) QQP QNLI SST-2 CoLA STS-B MRPC RTE Average
MEmeL 84.8/83.9 71.7 90.9 93.6 54.5 86.6 87.6 66.4 80.0
MEmeL-T 84.9/83.7 71.5 90.5 93.9 52.1 85.9 88.7 68.1 79.9
Table 6. Validation results on ImageNet-1k dataset. See Table. 2 for detailed settings. The metric is the higher the better. The best results are indicated in bold font.
Model Swin-T Swin-S Swin-B
Size 28M 50M 88M
MEmeL 81.1 83.4 85.1
MEmeL-T 81.1 83.5 85.1
Table 7. Results on HKO-7 dataset. The settings are the same with Table. 2. The metrics are the smaller the better with the best results in bold fonts.
Model B-MAE B-MSE
MEmeL 1003.25 309.94
MEmeL-T 1002.46 309.52
@@ F.1. The Impact of Removing MEmeL During Test
The results are displayed in Table. 5, Table. 6, and Table. 7 for GLUE, ImageNet, and HKO-7 datasets, respectively. The best results are shown in bold fonts. The performance differences for the two settings are very close on all three tasks, indicating a stable performance when tested without the induction of MEmeL. Interestingly, on GLUE datasets, the performance when using MEmeL during testing is much more fluctuated. It would be valuable to study the dynamics of MEmeL in more depth as future work.
@@ G.3. Difference from Traditional Normalizations
Previous normalizations have focused on the interaction between neurons in a layer, whereas our metric is based on the behavior of each individual neuron. For example, a neuron could be activated at 0 and deactivated at 2, which would be considered outstanding for normalization, as it consistently outputs a large value. One can better understand the importance of its 0-value through monosemantic analysis.
G.4. On the Selection of Benchmark and Experimental Settings
Currently, middle- to large-scale models typically follow a pretrain-finetune training procedure, such as Bert. Pretraining is usually conducted on multiple datasets to learn general abilities like feature extraction. After pretraining, the model is finetuned for a specific task, which may sacrifice its general ability in favor of better performance.
@@ G.4. On the Selection of Benchmark and Experimental Settings
However, validating the idea and performing pretraining is much more computationally expensive compared to finetuning, especially for large models. For instance, using the naive method to pretrain LLAMA2-7B based on our A800 is estimated to cost over \$$450k$ and over 5800 GPU-days (Touvron et al., 2023) with our preliminary test. Emergence occurs primarily near $10^{22}$ FLOPs and costs more than 10500 GPU-days (Wei et al., 2022). Conducting corresponding validation experiments is impossible for most researchers without the assistance of top companies.
@@
G.5. Potential Directions of Emergence Learning
@@
G.6. Future Work of Partially Decreasing Monosemanticity

### v3

@@ E.1. Configuration of Base Models
For the language task, we insert 4 MEmeL layers after the 4 middle Attention layers of BERT-base (12 in total). Since MEmeL does not introduce any additional parameters, we can directly use their checkpoint after pretraining to finetune. The finetuning setting is the same as BERT for fairness, where each task is trained with a batch size of 32 and 3 epochs over the data for all GLUE tasks. For each task, we selected the best fine-tuning learning rate ($l\in\{5e-5,4e-5,3e-5,2e-5\}$) on the validation dataset (DEV). Baseline results are from Table 1 in BERT (Devlin et al., 2019).
For the image classification task, we also insert 4 MEmeL layers after the 4 middle Attention layers for each model. The training is conducted on ImageNet-1k for 30 epochs using pretrained checkpoints on ImageNet-22k. The settings are directly adopted from their scripts, except for the number of GPU cards. Baseline results are collected by running checkpoints from the github of Swin-Transformer with some updates.
For the precipitation task, we insert 6 MEmeL lay- ers after the hidden states of each ConvGRU module. The training is conducted on the HKO-7 dataset (Shi et al., 2017) of radar images with a granularity of $120\times 120$, which are centered in Hong Kong. These radar images cover a timespan from 2009 to 2015 and are captured every 6 minutes. The original rainfall intensity is mapped to a scale of [0, 255], forming a regression task. The pretraining settings are the same as theirs, optimizing for 20k steps. Based on that, we conduct 2k more steps finetuning to obtain the final output.
@@ F.1. The Impact of Removing MEmeL During Test
The results are displayed in Table 6, Table 7, and Table 8 for GLUE, ImageNet, and HKO-7 datasets, respectively. The best results are shown in bold fonts. The performance differences for the two settings are very close on all three tasks, indicating a stable performance when tested without the induction of MEmeL. Interestingly, on GLUE datasets, the performance when using MEmeL during testing is much more fluctuated. It would be valuable to study the dynamics of MEmeL in more depth as future work.
Table 9. The paired t-test on all 3 datasets. “Mean-Original” and ‘Mean-Ours” refers to the mean scores of the two methods on each dataset. t-statistic¿0 for GLUE and ImageNet as their metrics are the larger the better. t-statistic¡0 for HKO-7 as its metric is the smaller the better.
Dataset GLUE ImageNet HKO-7
p-value 0.02 0.07 0.06
t-statistic 2.77 3.46 -2.38
Mean-Original 79.6 83.1 656.69
Mean-Ours 80.5 83.3 648.49
F.2. Statistical Significance
To statistically verify the improvement from our method, we conduct the paired t-test on all three datasets. The results are given in Table. 9. Each score is obtained from the results of the Original and MEmeL-Tune. All the datasets show a 90% confidence level (p-value ¡0.1) supporting the hypothesis that our performance significantly differs from the baseline.
@@ G.3. Difference from Traditional Normalizations
Here we demonstrate that our method is greatly different from normalization.
Normalization calculates the mean and variance of multiple neurons in a layer given 1 single input.
For our method, the mean and variance are computed for each single neuron with different inputs.
The underlying reason is that a neuron is considered activated when its value of 1 input deviates from the mean among all the inputs (Our points). While for normalization, it should be smoothed if it deviates from other neurons.
For example, let’s consider 5 neurons in a layer $(h_{1},h_{2},h_{3},h_{4},h_{5})$. Now assume that given the $i^{th}$ input $x^{(i)}$, the values of the 5 neurons are $(h_{1}^{(i)},h_{2}^{(i)},h_{3}^{(i)},h_{4}^{(i)},h_{5}^{(i)})=$$(-1,1,-1,0,-2)$.
For normalization:
Calculating the mean and standard deviation over all these 5 values, which are both scalar:
$\mu=\sum_{t=1}^{5}h_{t}^{(i)}/5=-0.6,$
$\sigma=\sqrt{(\sum_{t=1}^{5}(h_{t}^{(i)}-\mu)/5)}=1.0.$
Normalize the values of the 5 neurons:
$Norm(h_{1}^{(i)},h_{2}^{(i)},h_{3}^{(i)},h_{4}^{(i)},h_{5}^{(i)})$ $=((-1,1,-1,0,-2)-\mu)/\sigma$
$=(-0.4,1.7,-0.4,0.6,-1.4)$
For our deactivation:
We will first dynamically calculate the mean and the standard deviation for each neuron based on all the previous inputs $\{x^{(t)}\}_{t=1}^{i-1}$, which would both be 5 values:
historical mean: $(\mu_{1},\mu_{2},\mu_{3},\mu_{4},\mu_{5})=(3,0,-1,1,-1),$
and standard deviation: $(\sigma_{1},\sigma_{2},\sigma_{3},\sigma_{4},\sigma_{5})=(1,1,1,1,1).$
According to our metrics MS, neuron $h_{1}$ with value $h_{1}^{(i)}=-1$ and historical mean $\mu=3$ is the most activated. After our process, the output will become
$RD({h_{1}^{(i)}},h_{2}^{(i)},h_{3}^{(i)},h_{4}^{(i)},h_{5}^{(i)})$ $=((h_{1}^{(i)})^{\prime},h_{2}^{(i)},h_{3}^{(i)},h_{4}^{(i)},h_{5}^{(i)})$
$=(\mathbf{3},1,-1,0,-2).$
G.4. Correctness of Equation 5
Equation 5 draws some misunderstanding that it should be written as $f_{2}(\mathbf{z})$. However, $f_{2}$ should have both $\mathbf{x}$ and $\mathbf{z}$ as inputs when the model is not sequential. For example, in Figure 4(b), when focusing on $\mathbf{z}^{3}$, $\mathbf{o}=f_{2}(\mathbf{z}^{3},\mathbf{x})$ which is not only a function of $\mathbf{z}^{3}$ but also $\mathbf{z}^{4}(\mathbf{x})$.
G.5. Discussion of Experimental Settings
Currently, middle- to large-scale models typically follow a pretrain-finetune training procedure, such as Bert. Pretraining is usually conducted on multiple datasets to learn general abilities like feature extraction. After that, the model is finetuned for a specific task, which may sacrifice its general ability for better performance.
@@ G.5. Discussion of Experimental Settings
However, validating the idea and performing pretraining is much more computationally expensive compared to finetuning, especially for large models. For instance, using the naive method to pretrain LLAMA2-7B based on our A800 is estimated to cost over \$$450k$ and over 5800 GPU-days (AI, 2023) with our preliminary test. Emergence occurs primarily near $10^{22}$ FLOPs and costs more than 10500 GPU-days (Wei et al., 2022). Conducting corresponding validation experiments is impossible for most researchers without the help of top companies.
@@
G.6. Potential Directions of Emergence Learning
@@
G.7. Partially Decreasing Monosemanticity